The age of Keith is 1:6 of Nick's age. 8 years later, the ratio of Keith's age to Nick's age is 1 : 2. How old is Keith now?
1
Expert's answer
2013-04-25T10:15:19-0400
Let's denote the Keith's age by K and the Nick's age by N. Then we've got a system of equations:

K = 6N, (1)
K + 8 = 2(N + 8). (2)

Let's solve it. Substituting (1) into (2) we obtain:

6N + 8 = 2(N + 8) ==>

6N + 8 = 2N + 16 ==>

4N = 8 ==> N = 4.

So, Keith's age is K = 6N = 6*4 = 24 years and Nick's age is N = 4 years.